OpenStudy (anonymous):

4x squared - 35x + 49

OpenStudy (anonymous):

You want to solve t?

OpenStudy (anonymous):

If so 4x^2-35x+49=0 4x^2-28x-7x+49=0 4x(x-7)-7(x-7)=0 (4x-7)(x-7)=0 x=7/4 or x=7

jhonyy9 (jhonyy9):

4x^2 -35x +49=0 x_1,_2=(35+/- sqrt(1225-784))/8 =(35+/- sqrt441)/8 =(35+/-21)/8 = 14/8 = 7/4 and 56/8 =7 x_1=7/4 x_2=7